# VideoUFO: A Million-Scale User-Focused Dataset for Text-to-Video Generation

### v1

@@ Abstract
Text-to-video generative models convert textual prompts into dynamic visual content, offering wide-ranging applications in film production, gaming, and education. However, their real-world performance often falls short of user expectations. One key reason is that these models have not been trained on videos related to some topics users want to create. In this paper, we propose VideoUFO, the first Video dataset specifically curated to align with Users’ FOcus in real-world scenarios. Beyond this, our VideoUFO also features: (1) minimal ($0.29\%$) overlap with existing video datasets, and (2) videos searched exclusively via YouTube’s official API under the Creative Commons license. These two attributes provide future researchers with greater freedom to broaden their training sources. The VideoUFO comprises over $1.09$ million video clips, each paired with both a brief and a detailed caption (description). Specifically, through clustering, we first identify $1,291$ user-focused topics from the million-scale real text-to-video prompt dataset, VidProM. Then, we use these topics to retrieve videos from YouTube, split the retrieved videos into clips, and generate both brief and detailed captions for each clip. After verifying the clips with specified topics, we are left with about $1.09$ million video clips. Our experiments reveal that (1) current $16$ text-to-video models do not achieve consistent performance across all user-focused topics; and (2) a simple model trained on VideoUFO outperforms others on worst-performing topics. The dataset is publicly available here under the CC BY 4.0 License.
footnotetext: *Corresponding author.
Figure 1: The glowing firefly: (a) generated by Sora [2] and (b) captured in a real video. The generated firefly is noticeably different from its real-life counterpart and thus unsatisfying. We attribute this primarily to a lack of exposure to such topics.
@@ 1 Introduction
Text-to-video generation aims to convert textual descriptions into dynamic visual content. Its applications are extensive and transformative, covering areas from creative media [4] and entertainment [43] to practical domains such as education [6], advertising [46], and assistance [14].
@@
Despite their popularity and usefulness, current text-to-video models often fail to meet users’ expectations in real-world applications. For example, when we ask Sora [2] to generate a video using the prompt “A firefly is glowing on a grass’s leaf on a serene summer night”, it fails to capture the concept of a glowing firefly while successfully generating grass and a summer night, as shown in Fig. 1 (a). From the data perspective, we infer this is mainly because Sora [2] has not been trained on firefly-related topics, while it has been trained on grass and night. Furthermore, if Sora [2] has seen the video shown in Fig. 1 (b), it will understand what a glowing firefly should look like.
@@
To this end, this paper presents VideoUFO, the first Video dataset curated specifically based on real-world Users’ FOcus in text-to-video generation. Such a dataset can help improve the alignment between text-to-video models and actual user needs. Specifically, we focus on: (1) curating VideoUFO by analyzing real users’ interests and scraping relevant videos; (2) comparing the VideoUFO with recent video datasets to highlight differences; and (3) demonstrating how our VideoUFO benefits video generation.
@@
The first dataset that aligns with real-world users’ focus in text-to-video generation. Our key idea is to analyze user-focused topics from user-provided prompts and then search for videos related to these topics. As shown in Fig. 2, the resulted VideoUFO comprises more than $\mathbf{1.09}$ million video clips spanning $\mathbf{1,291}$ user-focused topics. Specifically, we initiate by analyzing user-focused topics, which involves embedding all $1.67$ million user-provided prompts from VidProM [35], clustering these embeddings using K-means, and generating a topic for each cluster with GPT-4o [21], followed by combining similar topics. After obtaining these topics, we (1) search for these topics on YouTube, (2) segment the fetched videos into multiple semantically consistent short clips, (3) generate both brief and detailed captions for each clip, (4) filter out clips that do not contain the specific topics, and (5) assign each clip video quality scores that align with human perception. Note that although our current VideoUFO comprises about one million videos, it can be easily scaled up to ten million or more by sourcing additional videos for each topic. The VideoUFO can also be easily extended to the image-to-video domain by using the text and image prompts in TIP-I2V [34].
@@
Differences between VideoUFO and other recent video datasets. Recently, several video datasets have been released, including OpenVid-1M [20], HD-VILA-100M [41], InternVid [38], Koala-36M [33], LVD-2M [40], MiraData [9], Panda-70M [5], VidGen-1M [29], and WebVid-10M [3]. While inheriting their excellent attributes – such as large scale, accurate captions, and high resolution – our VideoUFO explores some novel directions: (1) Guided by real user focus. Specifically, whereas recent datasets are typically gathered from open-domain sources, we concentrate on topics that are focused by text-to-video users. This feature enables text-to-video models trained on our dataset to better cater to users, while avoiding unnecessary expansion of the dataset and wasting resources. (2) Introducing new data. Although these recent papers claim to contribute new datasets, most of them primarily introduce a new data pipeline – that is, they reprocess HD-VILA-100M [41] and are subsets of it. While their contributions are useful and meaningful, in theory, a generative model that has already been fully fitted on HD-VILA-100M [41] would not gain new video knowledge from them. In contrast, we collect new data from YouTube, with only $0.29\%$ of the videos overlapping with existing datasets. (3) Data compliance. When curating VideoUFO, we retrieve videos using YouTube’s official API and select only those with a Creative Commons license. In contrast, most recent datasets do not explicitly address the regulatory compliance of their data collection process. This feature grants researchers greater flexibility in using our data.
@@
    We present VideoUFO, the first video dataset curated based on the focus of real text-to-video users. This dataset comprises over $1.09$ million video clips spanning $1,291$ user-focused topics.
@@ 2 Related Works
Text-to-Video Generation. The introduction of Sora [2] has sparked significant research interest in text-to-video generation. Commercial models such as Movie Gen [23], Veo [27], Kling [1], PixelDance [13], and Seaweed [13] demonstrate strong performance and are increasingly being integrated into various industries. Meanwhile, open-source models like HunyuanVideo [12], LTX-Video [15], CogVideoX [42], Mochi-1 [30], and Pyramidal [8] empower researchers to experiment, customize, and enhance existing frameworks. Although they are powerful, their training data determines the upper limit of the video quality they can generate. In this paper, we propose a dataset, VideoUFO, which has the potential to help these models better cater to users’ preferences in real-world applications.
Text-Video Datasets. A text-video dataset consists of video clips paired with corresponding textual descriptions or captions. Many text-video datasets, such as OpenVid-1M [20], HD-VILA-100M [41], InternVid [38], Koala-36M [33], LVD-2M [40], MiraData [9], Panda-70M [5], VidGen-1M [29], and WebVid-10M [3], have been proposed to advance multimodal understanding and generation. However, these datasets are not designed to align with the preferences of text-to-video users, creating a gap between training data topics and real-world applications. In contrast, our dataset VideoUFO is collected from a user-focus perspective. To better cater users, future researchers can fine-tune text-to-video models with the VideoUFO or integrate it with existing datasets to develop new models.
Preference Alignment in Diffusion Models. While diffusion models have demonstrated impressive capabilities, there is an increasing demand to ensure their outputs align with human preferences. Direct Preference Optimization (DPO) [25], originally proposed to align large language models (LLMs) with human preferences, can also be effectively applied to diffusion models. For example, DiffusionDPO [31] fine-tunes text-to-image models using human comparison data to improve image generation quality. Additionally, VideoDPO [17] introduces a comprehensive preference scoring system that evaluates both visual quality and semantic alignment to enhance text-to-video generation. Unlike other approaches that align text-to-image or text-to-video models with attributes such as aesthetics, motion, consistency, and visual appeal, our goal is to improve the performance of text-to-video models on real-world users’ focused topics.
[…]
Figure 4: Each data point in our VideoUFO includes a video clip, an ID, a topic, start and end times, a brief caption, and a detailed caption. Beyond that, we evaluate each clip with six different video quality scores from VBench [7].
@@ 3 Curating VideoUFO
Analyzing text-to-video users’ focused topics. We aim to analyze the topics that real users focus on or prefer when generating videos from texts. Here, we use VidProM [35] as it is the only publicly available, million-scale text-to-video prompt dataset written by real users. First, we embed all $1.67$ million prompts from VidProM into $384$-dimensional vectors using $\mathtt{SentenceTransformers}$ [26]. Next, we cluster these vectors with K-means. Note that here we pre-set the number of clusters to a relatively large value, i.e., $2,000$, and merge similar clusters in the next step. Finally, for each cluster, we ask GPT-4o [21] to conclude a topic ($1\sim 2$ words). The prompt is shown in the Supplementary (Section 8). After merging singular and plural forms, restoring verbs to their base forms, removing duplicates, and manually verifying the topics – removing overly broad ones such as ‘animation’, ‘scene’, ‘movement’, and ‘film’ – we finally obtained $\mathbf{1,291}$ topics. The semantic distribution of these topics is shown in Fig. 3, and the full list is available in the Supplementary (Section 12).
@@
Collecting videos from YouTube. For each topic obtained, we use the official YouTube API to search for videos. Specifically, for each topic, we search for approximately $500$ videos by relevance, with each video meeting the following criteria: (1) it is shorter than $4$ minutes, (2) it has a resolution of $720$p or higher, and (3) it is licensed under Creative Commons. These requirements ensure that the collected videos are suitable and freely usable for training video generation models. In the end, we obtain $586,490$ videos from YouTube. We compare the YouTube IDs of these videos with those in existing datasets, including OpenVid-1M [20], HD-VILA-100M [41], InternVid [38], Koala-36M [33], LVD-2M [40], MiraData [9], Panda-70M [5], VidGen-1M [29], and WebVid-10M [3], and find that only $1,675$ IDs ($0.29\%$) are already present in these datasets. This suggests that our VideoUFO introduces novel information or knowledge, which can serve to expand the range of existing training sources.
@@
Splitting videos and generating captions. After obtaining the videos, we segment them into multiple semantics-consistent short clips following the steps in curating Panda-70M [5]. The process includes shot boundary detection, stitching, and video splitting, producing a total of $\mathbf{3,181,873}$ clips. To facilitate the training of text-to-video models on our VideoUFO, we generate both brief and detailed captions for each video clip. For the brief captions, we utilize the video captioning model provided by Panda-70M [5]. For the detailed captions, we adopt the pipeline used in $\mathtt{Open}$-$\mathtt{Sora}$-$\mathtt{Plan}$ v1.3.0 [16], which employs $\mathtt{QWen2}$-$\mathtt{VL}$-$\mathtt{7B}$ [32] for video annotation. In Fig. 5 (a) and (b), we present the statistical distribution of the lengths of brief and detailed captions, respectively.
@@
Verifying clips. We notice that, although the videos are searched by topics, not every clip contains our intended topics. A straightforward solution is to use GPT-4o [21] to verify whether a given clip corresponds to a specific topic. However, while effective, this approach would be prohibitively expensive for verifying more than $3$ million video clips. As an alternative, we use the detailed caption (instead of the original clip) and feed it into GPT-4o mini (instead of GPT-4o) to verify whether it belongs to a specific topic. Since (1) GPT-4o mini and GPT-4o have similar capabilities in basic language understanding, and (2) the video understanding model provides detailed descriptions of the videos, we effectively complete the final verification step at a significantly lower cost. After verifying, there are $\mathbf{1,091,712}$ remaining clips. In Fig. 5 (c), we present the statistical distribution of clips duration.
Video quality assessment. To further support research in text-to-video generation, we evaluate the quality of each clip in VideoUFO. Specifically, we adopt six different video quality assessment metrics from VBench [7], which automatically assess videos and align well with human perception. The evaluation dimensions are: (1) subject consistency: assesses whether the main subject’s identity and appearance remain consistent throughout the video; (2) background consistency: evaluates the temporal stability and uniformity of the video’s background; (3) motion smoothness: measures the continuity and fluidity of movements within the video; (4) dynamic degree: assesses the level of activity and variation in motion present in the video; (5) aesthetic quality: judges the visual appeal and attractiveness of the video; and (6) imaging quality: examines the clarity, brightness, and color accuracy of the video. The statistical distributions of scores assessed by these six metrics are shown in the Supplementary (Fig. 9).
Extension. Future researchers can easily extend our VideoUFO in three aspects: (1) Scaling up. Although our VideoUFO has already reached a million-scale level, future researchers may still want to scale it up to 10 million or even 100 million. They can easily achieve this by leveraging our extracted topics to search more videos on platforms such as YouTube and TikTok. (2) New focusing. When we curate VideoUFO, the only available text-to-video prompt dataset is VidProM [35], which is used to analyze user focus and preferences. In the future, user focus may change, and other text-to-video prompt datasets may emerge for analyzing it. Future researchers can easily use our topic extraction pipeline to study these new focuses. (3) Image-to-video. Our current focus is text-to-video, which is the most common approach in the video generation community. Meanwhile, we also notice that image-to-video is gaining popularity. Future researchers can use text and image prompts, such as those in TIP-I2V [34], to analyze the focus of image-to-video users and collect corresponding datasets.
@@
Table 1: The comparison between the proposed VideoUFO with other recent video datasets based on their fundamental attributes. Unlike previous collections, our VideoUFO is derived directly from real user-focused topics and also offers a more flexible license, while remaining comparable to these datasets in other key aspects. “##\##” and “$\ \bar{}$ ” are abbreviations for “numbers” and “average”, respectively.
  Dataset
Venue
Caption
##\##Vid. $\bar{}$ $\bar{}$ Resolution
@@
WebVid-10M [3] ICCV21 Download 10M 18.0s 14.2 $<$360p Open 1,000 Retracted
HD-VILA-100M [41] CVPR22 Recognition 103M 13.4s 32.5 720p Open 648 R-UDA
InternVid [5] ICLR24 Generated 234M 11.7s 17.6 720p Open 1,051 Apache 2.0
Panda-70M [5] CVPR24 Generated 70M 8.5s 13.2 720p Open 719 R-UDA
LVD-2M [40] NeurIPS24 Generated 2M 20.2s 88.7 Diverse Open 814 R-UDA
MiraData [9] NeurIPS24 Generated 0.33M 72.1s 318.0 720p Open 639 GPL 3.0
Koala-36M [33] Arxiv24 Generated 36M 17.2s 202.1 720p Open 724 R-UDA
VidGen-1M [29] Arxiv24 Generated 1M 10.6s 89.3 720p Open 835 R-UDA
OpenVid-1M [20] ICLR25 Generated 1M 7.2s 127.3 Diverse Open 671 R-UDA
VideoUFO Ours Generated 1M 12.6s 155.5 720p Users $\mathbf{1,291}$ CC BY

### v2

@@ Abstract
Text-to-video generative models convert textual prompts into dynamic visual content, offering wide-ranging applications in film production, gaming, and education. However, their real-world performance often falls short of user expectations. One key reason is that these models have not been trained on videos related to some topics users want to create. In this paper, we propose VideoUFO, the first Video dataset specifically curated to align with Users’ FOcus in real-world scenarios. Beyond this, our VideoUFO also features: (1) minimal ($0.29\%$) overlap with existing video datasets, and (2) videos searched exclusively via YouTube’s official API under the Creative Commons license. These two attributes provide future researchers with greater freedom to broaden their training sources. The VideoUFO comprises over $1.09$ million video clips, each paired with both a brief and a detailed caption (description). Specifically, through clustering, we first identify $1,291$ user-focused topics from the million-scale real text-to-video prompt dataset, VidProM. Then, we use these topics to retrieve videos from YouTube, split the retrieved videos into clips, and generate both brief and detailed captions for each clip. After verifying the clips with specified topics, we are left with about $1.09$ million video clips. Our experiments reveal that (1) current $16$ text-to-video models do not achieve consistent performance across all user-focused topics; and (2) a simple model trained on VideoUFO outperforms others on worst-performing topics. The dataset and code are publicly available here and here under the CC BY 4.0 License.
@@ 1 Introduction
Text-to-video generation aims to convert textual descriptions into dynamic visual content. Its applications are extensive and transformative, covering areas from creative media bender2023coexistence and entertainment yu2024barriers to practical domains such as education garcia2024pre , advertising zheng2024open , and assistance li2024endora .
Figure 2: The glowing firefly: (a) generated by Sora openai2024sora and (b) captured in a real video. The generated firefly is noticeably different from its real-life counterpart and thus unsatisfying. We attribute this primarily to a lack of exposure to such topics.
@@
Despite their popularity and usefulness, current text-to-video models often fail to meet users’ expectations in real-world applications. For example, when we ask Sora openai2024sora to generate a video using the prompt “A firefly is glowing on a grass’s leaf on a serene summer night”, it fails to capture the concept of a glowing firefly while successfully generating grass and a summer night, as shown in Fig. 2 (a). From the data perspective, we infer this is mainly because Sora openai2024sora has not been trained on firefly-related topics, while it has been trained on grass and night. Furthermore, if Sora openai2024sora has seen the video shown in Fig. 2 (b), it will understand what a glowing firefly should look like.
@@
To this end, this paper presents VideoUFO, the first Video dataset curated specifically based on real-world Users’ FOcus in text-to-video generation. Such a dataset can help improve the alignment between text-to-video models and actual user needs. Specifically, we focus on: (1) curating VideoUFO by analyzing real users’ interests and scraping relevant videos; (2) comparing VideoUFO with other datasets to highlight differences; and (3) showing how VideoUFO benefits video generation.
@@
The first dataset that aligns with real-world users’ focus in text-to-video generation. Our key idea is to analyze user-focused topics from user-provided prompts and then search for videos related to these topics. As shown in Fig. 1, the resulted VideoUFO comprises more than $\mathbf{1.09}$ million video clips spanning $\mathbf{1,291}$ user-focused topics. Specifically, we initiate by analyzing user-focused topics, which involves embedding all $1.67$ million user-provided prompts from VidProM wang2024vidprom , clustering these embeddings using K-means, and generating a topic for each cluster with GPT-4o openai2024hello , followed by combining similar topics. After obtaining these topics, we (1) search for these topics on YouTube, (2) segment the fetched videos into multiple semantically consistent short clips, (3) generate both brief and detailed captions for each clip, (4) filter out clips that do not contain the specific topics, and (5) assign each clip video quality scores that align with human perception. Note that although our current VideoUFO comprises about one million videos, it can be easily scaled up to ten million or more by sourcing additional videos for each topic. The VideoUFO can also be easily extended to the image-to-video domain by using the text and image prompts in TIP-I2V wang2024tip .
@@
Differences between VideoUFO and other recent video datasets. Recently, several video datasets have been released, including OpenVid-1M nan2024openvid , HD-VILA-100M xue2022advancing , InternVid wanginternvid , Koala-36M wang2024koala , LVD-2M xiong2024lvd , MiraData ju2024miradata , Panda-70M chen2024panda , VidGen-1M tan2024vidgen , and WebVid-10M bain2021frozen . While inheriting their excellent attributes – such as large scale, accurate captions, and high resolution – our VideoUFO explores some novel directions: (1) Guided by real user focus. Specifically, whereas recent datasets are typically gathered from open-domain sources, we concentrate on topics that are focused by text-to-video users. This feature enables text-to-video models trained on our dataset to better cater to users, while avoiding unnecessary expansion of the dataset and wasting resources. (2) Introducing new data. Although these recent papers claim to contribute new datasets, most of them primarily introduce a new data pipeline – that is, they reprocess HD-VILA-100M xue2022advancing and are subsets of it. While their contributions are useful and meaningful, in theory, a generative model that has already been fully fitted on HD-VILA-100M xue2022advancing would not gain new video knowledge from them. In contrast, we collect new data from YouTube, with only $0.29\%$ of the videos overlapping with existing datasets. (3) Data compliance. When curating VideoUFO, we retrieve videos using YouTube’s official API and select only those with a Creative Commons license. In contrast, most recent datasets do not explicitly address the regulatory compliance of their data collection process. This feature grants researchers greater flexibility in using our data.
@@
    We present VideoUFO, the first video dataset curated based on the focus of real text-to-video users. This dataset comprises over $1.09$ million clips spanning $1,291$ user-focused topics.
@@ 2 Related Works
Text-to-Video Generation. The introduction of Sora openai2024sora has sparked significant research interest in text-to-video generation. Commercial models such as Movie Gen polyak2024moviegen , Veo veo2024 , Kling kuaishou_kling , PixelDance bytedance2024pixeldance_seaweed , and Seaweed bytedance2024pixeldance_seaweed demonstrate strong performance and are increasingly being integrated into various industries. Meanwhile, open-source models like HunyuanVideo kong2024hunyuanvideo , LTX-Video ltxvideo , CogVideoX yang2024cogvideox , Mochi-1 genmo2024mochi , and Pyramidal jin2024pyramidal empower researchers to experiment, customize, and enhance existing frameworks. Although they are powerful, their training data determines the upper limit of generated videos’ quality. This paper proposes a dataset, VideoUFO, which has the potential to help these models better cater to users’ preferences in real-world applications.
Figure 3: Each data point in our VideoUFO includes a video clip, an ID, a topic, start and end times, a brief caption, and a detailed caption. Beyond that, we evaluate each clip with six different video quality scores from VBench huang2024vbench .
Text-Video Datasets. A text-video dataset consists of video clips paired with corresponding textual descriptions or captions. Many text-video datasets, such as OpenVid-1M nan2024openvid , HD-VILA-100M xue2022advancing , InternVid wanginternvid , Koala-36M wang2024koala , LVD-2M xiong2024lvd , MiraData ju2024miradata , Panda-70M chen2024panda , VidGen-1M tan2024vidgen , and WebVid-10M bain2021frozen , have been proposed to advance multimodal understanding and generation. However, these datasets are not designed to align with the preferences of text-to-video users, creating a gap between training data topics and real-world applications. In contrast, our dataset VideoUFO is collected from a user-focus perspective. To better cater users, future researchers can fine-tune text-to-video models with the VideoUFO or integrate it with existing datasets to develop new models.
Preference Alignment in Diffusion Models. While diffusion models have demonstrated impressive capabilities, there is an increasing demand to ensure their outputs align with human preferences. Direct Preference Optimization (DPO) rafailov2023direct , originally proposed to align large language models (LLMs) with human preferences, can also be effectively applied to diffusion models. For example, DiffusionDPO wallace2024diffusion fine-tunes text-to-image models using human comparison data to improve image generation quality. Additionally, VideoDPO liu2024videodpo introduces a comprehensive preference scoring system that evaluates both visual quality and semantic alignment to enhance text-to-video generation. Unlike other approaches that align text-to-image or text-to-video models with attributes such as aesthetics, motion, consistency, and visual appeal, our goal is to improve the performance of text-to-video models on real-world users’ focused topics.
@@ 3 Curating VideoUFO
Analyzing text-to-video users’ focused topics. We aim to analyze the topics that real users focus on or prefer when generating videos from texts. Here, we use VidProM wang2024vidprom as it is the only publicly available, million-scale text-to-video prompt dataset written by real users. First, we embed all $1.67$ million prompts from VidProM into $384$-dimensional vectors using $\mathtt{SentenceTransformers}$ reimers2019sentencebert . Next, we cluster these vectors with K-means. Note that here we pre-set the number of clusters to a relatively large value, i.e., $2,000$, and merge similar clusters in the next step. Finally, for each cluster, we ask GPT-4o openai2024hello to conclude a topic ($1\sim 2$ words). The prompt is shown in the Appendix (Section A). After merging singular and plural forms, restoring verbs to their base forms, removing duplicates, and manually verifying the topics – removing overly broad ones such as ‘animation’, ‘scene’, ‘movement’, and ‘film’ – we finally obtained $\mathbf{1,291}$ topics. The semantic distribution of these topics is shown in Fig. 4.
@@
Collecting videos from YouTube. For each topic obtained, we use the official YouTube API to search for videos. Specifically, for each topic, we search for approximately $500$ videos by relevance, with each video meeting the following criteria: (1) it is shorter than $4$ minutes, (2) it has a resolution of $720$p or higher, and (3) it is licensed under Creative Commons. These requirements ensure that the collected videos are suitable and freely usable for training video generation models. In the end, we obtain $586,490$ videos from YouTube. We compare the YouTube IDs of these videos with those in existing datasets, including OpenVid-1M nan2024openvid , HD-VILA-100M xue2022advancing , InternVid wanginternvid , Koala-36M wang2024koala , LVD-2M xiong2024lvd , MiraData ju2024miradata , Panda-70M chen2024panda , VidGen-1M tan2024vidgen , and WebVid-10M bain2021frozen , and find that only $1,675$ IDs ($0.29\%$) are already present in these datasets. This suggests that our VideoUFO introduces novel information or knowledge, which can serve to expand the range of existing training sources.
@@
Splitting videos and generating captions. After obtaining the videos, we segment them into multiple semantics-consistent short clips following the steps in curating Panda-70M chen2024panda . The process includes shot boundary detection, stitching, and video splitting, producing a total of $\mathbf{3,181,873}$ clips. To facilitate the training of text-to-video models on our VideoUFO, we generate both brief and detailed captions for each video clip. For the brief captions, we utilize the video captioning model provided by Panda-70M chen2024panda . For the detailed captions, we adopt the pipeline used in $\mathtt{Open}$-$\mathtt{Sora}$-$\mathtt{Plan}$ v1.3.0 lin2024open , which employs $\mathtt{QWen2}$-$\mathtt{VL}$-$\mathtt{7B}$ Qwen2VL for video annotation. In the Appendix (Fig. 10 (a) and (b)), we present the statistical distribution of the lengths of brief and detailed captions, respectively.
@@
Verifying clips. We notice that, although the videos are searched by topics, not every clip contains our intended topics. A straightforward solution is to use GPT-4o openai2024hello to verify whether a given clip corresponds to a specific topic. However, while effective, this approach would be prohibitively expensive for verifying more than $3$ million video clips. As an alternative, we use the detailed caption (instead of the original clip) and feed it into GPT-4o mini (instead of GPT-4o) to verify whether it belongs to a specific topic. Since (1) GPT-4o mini and GPT-4o have similar capabilities in basic language understanding, and (2) the video understanding model provides detailed descriptions of the videos, we effectively complete the final verification step at a significantly lower cost. After verifying, there are $\mathbf{1,091,712}$ remaining clips. In the Appendix (Fig. 10 (c)), we present the statistical distribution of clips duration.
@@
Table 1: The comparison between the VideoUFO with recent video datasets based on their fundamental attributes. Unlike previous collections, our VideoUFO is derived directly from real user-focused topics and also offers a more flexible license, while remaining comparable to these datasets in other aspects. “##\##” and “$\ \bar{}$” are abbreviations for “numbers” and “average”, respectively.
  Dataset ##\##Vid. $\bar{}$ $\bar{}$ Resolution
@@
WebVid-10M bain2021frozen 10M 18.0s 14.2 $<$360p Open 1,000 Retracted
HD-VILA-100M xue2022advancing 103M 13.4s 32.5 720p Open 648 R-UDA
InternVid chen2024panda 234M 11.7s 17.6 720p Open 1,051 Apache 2.0
Panda-70M chen2024panda 70M 8.5s 13.2 720p Open 719 R-UDA
LVD-2M xiong2024lvd 2M 20.2s 88.7 Diverse Open 814 R-UDA
MiraData ju2024miradata 0.33M 72.1s 318.0 720p Open 639 GPL 3.0
Koala-36M wang2024koala 36M 17.2s 202.1 720p Open 724 R-UDA
VidGen-1M tan2024vidgen 1M 10.6s 89.3 720p Open 835 R-UDA
OpenVid-1M nan2024openvid 1M 7.2s 127.3 Diverse Open 671 R-UDA
VideoUFO 1M 12.6s 155.5 720p Users $\mathbf{1,291}$ CC BY
Video quality assessment. To further support research in text-to-video generation, we evaluate the quality of each clip in VideoUFO. Specifically, we adopt six different video quality assessment metrics from VBench huang2024vbench , which automatically assess videos and align well with human perception. The evaluation dimensions are: (1) subject consistency: assesses whether the main subject’s identity and appearance remain consistent throughout the video; (2) background consistency: evaluates the temporal stability and uniformity of the video’s background; (3) motion smoothness: measures the continuity and fluidity of movements within the video; (4) dynamic degree: assesses the level of activity and variation in motion present in the video; (5) aesthetic quality: judges the visual appeal and attractiveness of the video; and (6) imaging quality: examines the clarity, brightness, and color accuracy of the video. The statistical distributions of scores assessed by these six metrics are shown in the Appendix (Fig. 11).
Extension. Future researchers can easily extend our VideoUFO in three aspects: (1) Scaling up. Although our VideoUFO has already reached a million-scale level, future researchers may still want to scale it up to 10 million or even 100 million. They can easily achieve this by leveraging our extracted topics to search more videos on platforms such as YouTube and TikTok. (2) New focusing. When we curate VideoUFO, the only available text-to-video prompt dataset is VidProM wang2024vidprom , which is used to analyze user focus and preferences. In the future, user focus may change, and other text-to-video prompt datasets may emerge for analyzing it. Future researchers can easily use our topic extraction pipeline to study these new focuses. (3) Image-to-video. Our current focus is text-to-video, which is the most common approach in the video generation community. Meanwhile, we also notice that image-to-video is gaining popularity. Future researchers can use text and image prompts, such as those in TIP-I2V wang2024tip , to analyze the focus of image-to-video users and collect corresponding datasets.